Shift Summary 9-10-99 Day

Greg Bock (bock@fnal.gov)
Fri, 10 Sep 1999 15:26:46 -0500

Shift Summary for day shift, 9-10-99
On Shift: Arun,GregB (+Ed, Sydney,..)

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):
Accounting of lost time:

========================================================================
Runs ended this shift: 14491 (3M events to tape)

Tapes brought to FCC this shift:
QKB144,QKC408-410,QKK408-410,QKS408-410

========================================================================
Accesses during shift:
ET low voltage trips due to high temp. Sydney and Arun investigate.
Noticeably warmer in the ET rack vicinity.

========================================================================
Overview of shift: 2 minutes of beam at beginning.
E799 split tapes labelled
gas checked
access to investigate ET trip
shift ends with a few pulses of low intensity poorly
targetted beam
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

======================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
Keep an eye on hall temperature, and ET temperature.
Figure out how to work the noisebusters that Leo brought